[Brahms-l] Shift report 20040208 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 08 2004 - 23:35:56 EST
------------------------------------------------------------
Date: Sun Feb  8 23:35:56 2004, Shift: 16:00-24:00
Supervisor: IA, with: 
Log book#15 pages 51 - 51

Shift Summary:
------------------------------------------------------------
1. Quench Link Interlock problems for MCR. We had only 2.5 hours of beam.

2. Data runs: 10211, 10212, 10213, 10216
 The settings for the entire shift :
  FS at 2.3 degrees, A1723
  MRS at 45 degrees, A2255
 ~600k trig2 and ~340k trig3 taken.

3. No detector or magnet problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 8 23:36:08 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 08 2004 - 23:36:28 EST